### v1

# Realization of the Tellegen Effect in Resonant Optical Metasurfaces

Abstract
Abstract
The nonreciprocal magnetoelectric effect in isotropic Tellegen materials enables exotic phenomena such as axion-modified electrodynamics [1, 2, 3, 4, 5, 6, 7, 8] and fosters the development of magnet-free nonreciprocal media [9, 10, 11]. Although predicted over 75 years ago [12] and observed in only a handful of natural materials with very low strength [13], the strong optical Tellegen effect has remained experimentally elusive. Here, we report the first experimental demonstration of a resonant optical isotropic Tellegen effect in a metasurface, showcasing a response that is 100 times greater than that of any known natural material. This optical metasurface, consisting of randomly distributed cobalt-silicon nanoscatterers with strong shape anisotropy, utilizes spontaneous magnetization to achieve a robust Tellegen effect without the need for an external magnetic field. In addition to the Tellegen response, the metasurface exhibits both gyroelectric and gyromagnetic effects [14], contributing to nonreciprocal cross-polarized light reflection. We introduce a technique to independently extract the amplitudes of these three effects using conventional magneto-optical single-side-illumination measurements. The observation of the resonant Tellegen effects in the optical frequency range may lead to the experimental observation of axionic electrodynamics and compact bias-free nonreciprocal optical devices.
Magnetoelectric coupling is a defining characteristic of bianisotropic materials, where a magnetic field induces electric polarization and, conversely, an electric field induces magnetization. Bianisotropic phenomena are categorized into reciprocal (commonly known as chirality) and nonreciprocal (Tellegen) effects [15, 16]. Chirality, which arises in materials that exhibit reciprocity but lack parity, has been extensively studied in natural and engineered materials  [17, 18, 19, 20]. In contrast, the nonreciprocal magnetoelectric effect, named after Bernard Tellegen, who first proposed it in 1948 [12], is extremely rare as it requires the simultaneous breaking of both parity symmetry and reciprocity [13, 14].
The isotropic Tellegen effect is expected to have a significant impact on both fundamental and applied physics [21]. It should be clearly distinguished from its anisotropic counterpart, the magnetochiral effect [22, 23], as it necessitates a diagonal magnetoelectric tensor with a non-zero trace. Seminal works have demonstrated that the electrodynamic equations governing isotropic Tellegen materials take the same form as those describing a hypothetical axion medium within the framework of quantum field theory [1, 2]. The axion is an elementary particle theorized to address two major issues in modern physics: the strong charge conjugation and parity (CP) problem [24] and the existence of dark matter [25]. Tellegen-type nonreciprocal materials have attracted considerable interest as potential enablers of exotic concepts of axion electrodynamics in practical materials, such as dyon quasiparticles, anyon statistics, the Witten effect, and axionic polaritons [2, 26, 27, 28]. They may also play a crucial role in the search for axions [6, 7, 5]. Moreover, Tellegen materials promise a range of exotic phenomena, especially in the optical regime, including nonreciprocal light reflection [29, 10] and transmission [9], photonic topological edge states [30], PT-symmetric magnetoelectric energy density [31], and spin-dependent thermal radiation [32].
While the isotropic Tellegen effect could be measured at microwave frequencies for certain multiferroic materials [21, 33, 34], it already becomes two orders of magnitude smaller than the refractive index in the terahertz range [35] and five orders of magnitude smaller in the optical regime [36, 37]. This drastic reduction at optical frequencies is attributed to the fact that both cyclotron and Larmor frequencies of electrons responsible for breaking the Lorentz reciprocity typically lie in the gigahertz range. Thus, artificial Tellegen composites designs have been confined to the microwave frequencies: from individual constituents (meta-atoms) [38, 39, 40] to the two-dimensional arrays (metasurfaces) [41, 42, 11]. The quest for optical Tellegen effect has inspired three recent conceptual proposals for Tellegen metamaterials. The first two approaches, based on rapid temporal modulation of material properties [43, 44] and multilayer antiferromagnetic configuration [8, 45], are not achievable with the current manufacturing technology. The third theoretical proposal based on metasurfaces with spontaneous magnetization in magnetic two-component nanoparticles has not been experimentally validated either [10].
Here, for the first time, we fabricate and experimentally characterize an optical Tellegen metasurface with a diagonal magnetoelectric tensor. The metasurface consists of nanoscale cone-shaped scatterers, composed of amorphous silicon and cobalt. Due to strong shape anisotropy, the cobalt part is in a single-domain magnetic state at fabrication, ensuring that the metasurface displays the Tellegen effect even in the absence of external magnetization. We further demonstrate that the optical Tellegen metasurface also exhibits significant gyroelectric and nontrivial gyromagnetic [46] responses. As all three effects collectively contribute to the non-reciprocal reflection of the cross-polarized light from the metasurface, we propose an original method to independently determine their respective amplitudes using conventional magneto-optical measurements. The metasurface displays a Tellegen parameter on the order of $10^{-3}$, which aligns well with theoretical and simulation predictions and is at least 100 times larger than values observed in natural materials such as chromium oxide [47]. Moreover, the chosen nanoscatterers geometry enables producing large-scale optical Tellegen metasurfaces using high-throughput bottom-up nanofabrication, allowing straightforward scalability to hundreds of square centimeters.
Metasurface design
The isotropic Tellegen effect is primarily an interface phenomenon manifesting itself as a nonreciprocal cross-polarized light reflection at a vacuum-matter interface. For spatially uniform slabs, the effect remains undetectable in light transmission [29]. Unlike magneto-optical effects of both gyroelectric and gyromagnetic types, Tellegen media exhibit a direction-independent Kerr effect due to isotropy and do not exhibit the Faraday effect [10]. This distinction makes measuring the Tellegen effect more challenging, as its observable manifestation—the “isotropic” Kerr effect—depends solely on the concentration of Tellegen meta-atoms rather than their total number or the slab thickness [29, p. 261]. Therefore, the fabrication and experimental verification of high-density optical Tellegen metasurfaces, where interface phenomena are readily observable, represent the most straightforward approach to exploring the Tellegen effect.
Figure 1: Optical Tellegen metasurface design. (a) Illustration of the metasurface with optimized optical Tellegen meta-atoms consisting of cobalt (top) and silicon (bottom) parts. The metasurface is illuminated from opposite directions by normally incident light (a slight oblique incidence is depicted for visual clarity). The gradient color indicates that the reflected and transmitted light contain both co- and cross-polarized components. The inset depicts a close-up image of the meta-atom with close to single-domain vertical magnetization ${\bf M\mit}_{\rm s}$ in the cobalt. It also shows the high-frequency magnetic field distribution inside the nanocone when illuminated by a standing wave with an electric field antinode positioned at the center of the meta-atom. The field distribution is calculated at a free-space wavelength of 830 nm and reveals an $x$-oriented magnetic dipole moment ${\bf m\mit}$ induced solely by the incident $x$-polarized electric field through the isotropic (diagonal) Tellegen effect. (b)–(e) Simulated complex polarizabilities of the metasurface in (a), normalized by the free-space impedance $\eta_{0}$.
We select the Tellegen meta-atom as a nanoscale cone-shaped scatterer, shown in the inset of Fig. 1(a). The upper part of the nanocone is cobalt, with a height $h_{\rm Co}=119$ nm and a base diameter $D_{\rm Co}=76$ nm. The lower part is amorphous silicon, with a height $h_{\rm Si}=20$ nm and a base diameter $D_{\rm Si}=89$ nm. This geometry has an apex angle of approximately $35^{\circ}$, enabling straightforward fabrication via bottom-up hole-mask colloidal lithography (HCL). Cobalt possesses one of the largest magneto-optical parameters in the visible spectrum [48]. The optimal base diameter of the cobalt part, $D_{\rm Co}$, was determined by fabricating test metasurfaces consisting of cobalt-only nanocones to achieve a configuration closest to the ideal single-domain state (see Supplementary Section 1) that arises due to the strong shape anisotropy [49] and remains unaffected by the addition of the silicon. That is, the cobalt part can sustain strong vertical remanent magnetization close to the saturation magnetization $M_{\rm s}$ in the absence of an external magnetic field. The optimal height of the silicon part was determined by full-wave simulations to induce a magnetic-type Mie resonance [10], thereby enhancing the Tellegen response of the meta-atom (see Supplementary Section 2). The magnetic field distribution of this resonant mode inside the nanocone is depicted in the inset of Fig. 1(a) at a free-space wavelength of 830 nm. Here, the excitation consists of a standing wave with an $x$-polarized electric field antinode located at the center of the meta-atom. This field alone induces an $x$-oriented magnetic dipole moment ${\bf m\mit}$ inside the nanocone – a distinctive signature of the Tellegen effect [16].
The nanocone possesses the magnetic point group $C_{\infty v}(C_{\infty})$, indicating that it can support only gyroelectric, gyromagnetic, reciprocal bianisotropic (omega coupling), and nonreciprocal (diagonal) Tellegen effects. Here, we refer to gyroelectric and gyromagnetic effects as those responsible for Kerr and Faraday rotation in materials with off-diagonal permittivity and permeability tensors, respectively. Notably, the symmetry of the meta-atom inherently prohibits any chirality. A detailed analysis of the electromagnetic material tensors corresponding to this point group can be found in [10]. It is important to highlight that the chosen meta-atom exhibits both parity and reciprocity breaking, leading to the emergence of the Tellegen effect even in the absence of silicon Mie resonator part. However, adding it results in a two-fold enhancement of the Tellegen effect due to the resonance mechanism (see Supplementary Section 2).
We first consider the metasurface in free space with meta-atoms in a square periodic lattice (Fig. 1(a).) We analyze the metasurface response using full-wave simulations and extract the collective polarizabilities, which relate the induced electric and magnetic dipole moments ($\mathbf{p}$ and $\mathbf{m}$) to the incident electric and magnetic fields ($\mathbf{E}$ and $\mathbf{H}$[16]:
$\begin{array}[]{c}\displaystyle{\bf p\mit}=\left[\alpha_{\rm e}\overline{\overline{I}}+\alpha_{\rm ge}\overline{\overline{J}}\right]\cdot{\bf E\mit}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf H\mit},\vspace{6pt}\\ \displaystyle{\bf m\mit}=\left[\alpha_{\rm m}\overline{\overline{I}}+\alpha_{\rm gm}\overline{\overline{J}}\right]\cdot{\bf H\mit}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf E\mit},\end{array}$ (1)
where $\overline{\overline{I}}$ represents the two-dimensional identity matrix, and $\overline{\overline{J}}$ denotes the antisymmetric (skew-symmetric) two-dimensional matrix characterized by the nonzero components $J_{yx}=-J_{xy}=1$. Light is incident on the metasurface along the $z$-axis (normal to the surface), which is also the uniaxial direction of the metasurface. The collective polarizabilities $\alpha_{\rm e}$ and $\alpha_{\rm m}$ represent the ordinary reciprocal electric and magnetic responses, respectively. The terms $\alpha_{\rm ge}$ and $\alpha_{\rm gm}$ correspond to the nonreciprocal gyroelectric and gyromagnetic responses, while $\alpha_{\rm te}$ and $\alpha_{\rm om}$ describe the nonreciprocal Tellegen and reciprocal omega responses, respectively. All the polarizabilities are defined with respect to the base of the nanocone.
Figure 2: Extraction of the independent gyrotropic, gyromagnetic, and Tellegen effects in a general nonreciprocal metasurface using magneto-optical measurements. (a)–(c) Three equivalent metasurfaces with identical meta-atoms and the same average surface densities on top of alumina spacers of three different thicknesses $h_{1}=0$ nm, $h_{2}=60$ nm, and $h_{3}=120$ nm. The metasurface are illuminated from the top by a normally incident light, and the Kerr rotation $\theta$ and ellipticity $\epsilon$ are extracted (an oblique incidence is depicted only for visual clarity). (d) The complex amplitudes of the coefficients introduced in Eq. 3 plotted as functions of the free-space wavelength $\lambda$ and the alumina spacer thickness $h$. These coefficients quantify the contributions of the gyroelectric, gyromagnetic, and Tellegen effects to cross-polarized light reflection.
The Tellegen polarizability $\alpha_{\rm te}$ can be calculated for the considered metasurface as
$\alpha_{\rm te}=\frac{ia^{2}}{2\omega}(R_{\rm f}^{\rm cr}-R_{\rm b}^{\rm cr}),$ (2)
where $R_{\rm f}^{\rm cr}$ and $R_{\rm b}^{\rm cr}$ denote complex-amplitude cross-polarized reflection coefficients for the forward and backward illuminations, respectively, $a=320$ nm is the metasurface period, and $i$ is the imaginary unit (we adopt time-harmonic oscillations of the form $e^{i\omega t}$). This and other expressions for the unknown polarizabilities are provided in Supplementary Section 3. As observed in (2), extracting the Tellegen effect in the metasurface requires illumination from both sides. This is due to the fact that all three effects—gyroelectric, gyromagnetic, and Tellegen—are nonreciprocal and collectively contribute to the cross-polarized light reflection. Consequently, to determine separately the Tellegen effect in a metasurface that may also display gyroelectric and/or gyromagnetic responses, one must evaluate the specific linear combination of the two complex reflection coefficients. This necessitates precise measurement of both their complex amplitudes and phases.
By evaluating the full scattering matrix of the metasurface through full-wave simulations, we compute the three nonreciprocal collective polarizabilities (see Supplementary Section 3), as shown in Figs. 1(b)–(d). From Fig. 1(e), it can be observed that all three polarizabilities exhibit similar magnitudes, necessitating illumination from both sides for their accurate extraction. At approximately 830 nm, the Tellegen response exhibits a Lorentzian-type resonance. Additionally, the Tellegen meta-atom demonstrates a resonant gyromagnetic response, whose presence at optical frequencies was only recently observed in metamaterials [46].
Crucially, when the metasurface is experimentally realized on a substrate, measuring the complex-valued reflection coefficients for both illumination directions becomes exceedingly challenging. Due to the finite bandwidths of supercontinuum lasers required for broadband measurements, the large thickness of a typical substrate would exceed the coherence length, driving the system into an incoherent regime. This would make phase measurements of the reflection coefficients for both forward and backward illuminations in practice impossible. The substantial total thickness variation of the substrate, caused by non-uniform roughness, would further complicate phase measurements, as determining the thickness of such a thick substrate with high precision is challenging. Without phase information, the Tellegen response cannot be accurately extracted, as the gyroeletric and gyromagnetic effects would contribute to the measurements, as discussed above.
Experimental extraction of the Tellegen effect
To address these challenges, we introduce a novel method: we fabricate and characterize three Tellegen metasurfaces, measuring their reflection responses under illumination from a single side. These metasurfaces are structurally identical, each composed of the same type of meta-atoms arranged with an equal average surface density, as illustrated in Figs. 2(a)–(c). The key distinction lies in the thickness of the dielectric spacers on which they are mounted. Beneath each dielectric spacer is a continuous 100 nm-thick aluminum layer, followed by a glass ($\text{SiO}_{2}$) substrate with dimensions $24\,\text{mm}\times 24\,\text{mm}\times 0.4\,\text{mm}$, which ensures nearly complete reflection of incident light. The dielectric spacer is composed of alumina ($\text{Al}_{2}\text{O}_{3}$), chosen for its low optical loss. The thin and precisely controlled thickness of alumina spacer enables straightforward and reliable optical measurements of the complex reflection coefficient. As stated above, the cross-polarized reflection is influenced not only by the Tellegen effect but also by gyroelectric and gyromagnetic effects in the metasurface. To independently determine the amplitudes of these three effects, measurements from three metasurfaces are required.
For the metasurfaces considered in Figs. 2(a)–(c), the cross-polarized reflection of normally incident light linearly depends on the three polarizabilities according to (see Supplementary Section 4):
$\begin{array}[]{c}\displaystyle R_{\text{cr}}=C_{\rm ge}\,\eta_{0}\alpha_{\rm ge}+C_{\rm gm}\,\alpha_{\rm gm}/\eta_{0}+C_{\rm te}\alpha_{\rm te},\end{array}$ (3)
where we introduce $C_{\rm ge}$, $C_{\rm gm}$, and $C_{\rm te}$ as the weight amplitudes of the gyroelectric, gyromagnetic, and Tellegen effects in the cross-polarized reflected light. These amplitudes depend on the free-space wavelength $\lambda$, periodicity of the metasurface $a$, as well as material properties and thicknesses of the dielectric spacer and metallic layer. Importantly, these coefficients are independent of the meta-atoms and represent only the environment surrounding the metasurface. The free-space impedance $\eta_{0}$ is included in (3) to ensure that the polarizabilities remain normalized and have consistent dimensions.
The weight coefficients are plotted versus the free-space wavelength $\lambda$ and the dielectric spacer thickness $h$ in Fig. 2(d). For a given value of $\lambda$ the three weight coefficients exhibit maxima and minima along the $h$-axis, with the periodicity given by $\Delta h=\lambda/(2n_{\rm d})$ (see Supplementary Section 4), where $n_{\rm d}$ is the refractive index of the dielectric spacer. We see that for some specific values of the spacer thickness $h$, the gyroelectric (or gyromagnetic) effect is nearly the sole contributor to the cross-polarized reflection $R_{\mathrm{cr}}$. This implies that by selecting such spacer thicknesses, one can reliably isolate and measure the gyroelectric (or gyromagnetic) properties of the metasurface deposited on the spacer using a single measurement. This holds true regardless of the possible presence of other nonreciprocal effects. However, such independent extraction of gyroelectric (or gyromagnetic) effects is effective only at a single wavelength. To determine these effects over a broad wavelength range, measurements must still be performed on three metasurfaces with different spacer thicknesses. Further, a dielectric spacer at which $R_{\mathrm{cr}}$ depends solely on the Tellegen effect of the metasurface ($\alpha_{\rm te}$) even at a single wavelength is not achievable. As can be seen in Fig. 2(d), no spacer thickness $h$ exists for which both the gyroelectric and gyromagnetic contributions vanish simultaneously. Consequently, it is impossible to measure the Tellegen effect from a general nonreciprocal metasurface using a single metasurface and single-side illumination.
Generally, in our method of Tellegen effect extraction using three metasurfaces, one selects three arbitrary spacer thicknesses $h_{1}$, $h_{2}$, $h_{3}$ to obtain a system of three linearly independent equations for $R_{\rm cr}^{(1)}$, $R_{\rm cr}^{(2)}$, and $R_{\rm cr}^{(3)}$ as in (3), which can then be solved with respect to the unknown $\alpha_{\rm ge}$, $\alpha_{\rm gm}$, and $\alpha_{\rm te}$. The complex cross-polarized reflection coefficients are determined experimentally by performing conventional magneto-optical polar Kerr measurements [50, 51, 52]. Specifically, one can write for the $(j)$-th metasurface $R_{\rm cr}^{(j)}=(\theta_{j}+i\epsilon_{j})R_{\rm co}^{(j)}$, where $\theta$ and $\epsilon$ are the Kerr rotation and ellipticity, respectively, and $R_{\rm co}$ is the co-polarized complex reflection coefficient [14]. Without loss of generality, here we choose the dielectric spacer thicknesses as $h_{1}=0$ nm, $h_{2}=60$ nm, and $h_{3}=120$ nm approximately corresponding to the maxima of the three weight amplitudes within the wavelength range of interest (around 830 nm).
To determine the Tellegen effect of the proposed optical metasurface, we follow the schematics depicted in Fig. 2. First, we fabricate three samples with aforementioned spacer thicknesses (Figs. 2(a)–(c)). Using the magneto-optical measurements, we obtain three sets of the Kerr rotation and ellipticity $\theta_{j}$, $\epsilon_{j}$. Next, we find complex-valued reflection coefficients $R_{\rm cr}^{(j)}$ and weight amlpitudes $C_{\rm ge}$, $C_{\rm gm}$, and $C_{\rm te}$. Finally, by solving the system of three linear equations, we retrieve the unknown polarizabilities $\alpha_{\rm ge}$, $\alpha_{\rm gm}$, and $\alpha_{\rm te}$ (Supplementary Section 4).
Experimental realization of the optical Tellegen metasurface
Figure 3: Fabricated Tellegen metasurface and its characterization. (a)–(b) SEM images of one of the fabricated Tellegen metasurfaces with dielectric spacer thickness of $h=60$ nm (top and tilted views). The inset in (b) presents a zoomed-in image of a single meta-atom, artificially colored to visualize the cobalt and silicon parts with the averaged geometrical parameters of the nanocone used in full-wave simulations: $h_{\rm Si}=20$ nm, $h_{\rm Co}=104$ nm, $D_{\rm Si}=90$ nm, and $D_{\rm Co}=84$ nm. The averaged half-apex angle of the nanocone is $85^{\circ}$. (c) Schematic illustration of the polar magneto-optic Kerr effect (MOKE) experimental setup. Detailed information about the setup is available in the Methods section. (d)–(f) Experimental and simulated magneto-optical Kerr spectra for the three fabricated metasurfaces under an external magnetic field fixed at $B=1.3$ T. Here, the positive-value magnetic field denotes the direction from the base of the meta-atom to its apex. The simulated spectra correspond to periodic metasurfaces with geometric parameters averaged from the fabricated samples. (g)–(i) The gyroelectric, gyromagnetic, and Tellegen polarizabilities extracted from the data in (d)–(f).
Three metasurfaces were fabricated using hole-mask colloidal lithography (see Materials and Methods Section). Figures 3(a) and (b) present scanning electron microscopy (SEM) images of one of the fabricated metasurfaces. The images confirm the conical shape of the meta-atoms, according to the design, while meta-atoms distribution provides a quasi-uniform surface density. By extracting meta-atoms average surface density using the SEM images, we find the mean interparticle distance to be $a=295$ nm. The calculated mean base diameter of the meta-atoms is $D_{\rm Si}=90$ nm with a standard deviation of 23 nm. The average height of the cones $h_{\rm Co}+h_{\rm Si}$ is about 130 nm, which is slightly lower than the 140 nm expected for an idealized nanocone. However, the cones exhibit rounded apexes rather than sharp ones. This characteristic is also evident in the SEM images (inset in Fig. 3(b)) and AFM images (see Supplementary Section 5), where the apex appears circular rather than pointed. To ensure a fair comparison between experimental and simulated results, we modified the meta-atom topology in our full-wave simulations by replacing the idealized nanocone with a sharp apex with one featuring a rounded apex. An inset in Fig. 3(b) illustrates the geometric parameters of the simulated nanocone specified in the figure caption.
We measure the magneto-optical Kerr rotation angle ($\theta$) and Kerr ellipticity ($\varepsilon$) using a magneto-optical setup illustrated in Fig. 3(c) (see more details in Methods and Supplementary Section 6). The samples were positioned between the poles of an electromagnet generating a magnetic field of $B=\pm 1.3$ T, perpendicular to the measured surface. This field alternately reversed the spontaneous magnetization of the nanocones along their axes. The obtained hysteresis curves can be found in Supplementary Section 7. The incident light was directed normally onto the metasurface, passing through a hole in the electromagnet. In the measurements of the Tellegen effect, magneto-optical data are acquired under an external magnetic field of $B=1.3$ T. This ensures that the meta-atoms in all three metasurfaces remain in the same uniformly saturated magnetization state, making the metasurfaces magneto-optically equivalent and enabling the use of our extraction technique. Measuring the Tellegen effect in the absence of an external field is also possible but would require more costly and far less scalable nanofabrication that results in identical meta-atoms diameters across the metasurface.
Figure 4: Simulated results for purely Tellegen material systems. (a) A thin (50 nm) silicon nitride membrane with two identical Tellegen metasurfaces shown in Fig. 1(a) being deposited on its both sides. The membrane is illuminated from opposite directions by normally incident light (an oblique incidence is depicted for visual clarity). (b) Simulated complex collective polarizabilities of the membrane in (a), normalized by the free-space impedance $\eta_{0}$. (c) Kerr rotation angle and Kerr ellipticity for the membrane in (a) calculated for forward and backward light illuminations. (d) Isotropic Tellegen colloid consisting of randomly oriented nanocones inside water. The colloid exhibits the same magneto-optical Kerr effect regardless of the illumination direction. (e) Effective bulk Tellegen parameter of the colloid. (f) Kerr rotation angle and Kerr ellipticity for the colloid in (d).
Figures 3(d)–(f) present the experimental Kerr rotation and ellipticity for the three metasurfaces ($\theta_{1,2,3}$ and $\epsilon_{1,2,3}$) as functions of the free-space wavelength. For comparison, we also include the simulated Kerr parameters for periodic metasurfaces. It is important to note that these full-wave simulations cannot fully capture the complexity of the experimental metasurfaces, featuring meta-atoms with significant variations in their overall sizes (although with the same proportions) and quasi-random distances between neighboring meta-atoms. In the simulations, we assume a periodic arrangement of meta-atoms with an effective period of $a=330$ nm, using averaged geometric parameters of the meta-atoms depicted in the inset of Fig. 3b).
Despite the inherent differences between the quasi-random experimental metasurfaces and their periodic simulated counterparts, we observe both qualitative and satisfactory quantitative agreement between the measured and simulated Kerr spectra. All the spectra exhibit the characteristic Lorentzian dispersion. For the first metasurface with a spacer thickness of $h_{1}$, the simulated Kerr parameters are slightly larger in amplitude than the measured ones, although the spectral correspondence remains strong. For the second metasurface with thickness $h_{2}$, the experimental and simulated data show good agreement both spectrally and quantitatively. However, for the third metasurface with $h_{3}$, the simulated results are redshifted by approximately 150 nm relative to the experimental data.
By applying our proposed extraction technique (summarized in Fig. 2), we calculate the polarizabilities of the experimental metasurfaces using both measured and simulated data from Figs. 3(d)–(f). The extracted polarizability spectra are plotted in Figs. 3(g)–(i). One can observe a good qualitative agreement between the measured and simulated data, both spectrally and in terms of amplitude. All three effects—gyroelectric, gyromagnetic, and Tellegen—exhibit comparable amplitudes within the wavelength range of interest. Comparing these results with the simulated polarizabilities of the idealized meta-atom in Fig. 1, we observe similar response, although the resonance frequencies of the effects have shifted. These quantitative differences stem from the substantial geometric variations between the idealized meta-atom in Fig. 1(a) and the realistic meta-atom in Fig. 3(b). Furthermore, in our extraction, we model the array of meta-atoms as a homogenized sheet of electric and magnetic currents with deeply subwavelength thickness. This introduces an additional source of inaccuracy, as the actual meta-atom height is on the order of $\lambda/6$. Nevertheless, both the simulated and measured results provide strong evidence for the resonant nature of the diagonal Tellegen effect in the proposed optical metasurface. The magnitudes of the Tellegen polarizability $|\alpha_{\rm te}|$, extracted from both the measured data and the simulations for the idealized meta-atoms in Fig. 1 at 830 nm, reaches approximately $6\times 10^{-33}$ ${\rm m}^{2}\cdot{\rm s}$. These results demonstrate that the observed amplitude of the Tellegen effect is comparable to that of the resonant gyroelectric effect present in the metasurface. This implies that the Tellegen effect is sufficiently strong to be relevant for magneto-optical applications. In the following section, we calculate and discuss the magnitude of the observed value and determine the corresponding bulk magnetoelectric parameter.
Discussion
The demonstrated Tellegen metasurface can be further used for the implementation of purely Tellegen metamaterials. In this section, we numerically explore the physics and observable effects in two conceptual material systems that can be built based on the proposed Tellegen metasurfaces: a Tellegen membrane and an isotropic Tellegen colloid, depicted in Figs. 4(a) and (d), respectively. The former consists of two proposed metasurfaces deposited on the top and bottom sides of a nanoscale-thin silicon nitride membrane, designed to have spontaneous magnetizations in opposite directions, as shown in Fig. 4(a). While this configuration requires challenging fabrication steps and sophisticated magnetic field sequences, it can be realized over small footprints [53]. The second material system, an isotropic Tellegen colloid, consists of randomly oriented nanocones dispersed within a host medium such as water [10]. Both the membrane and colloid share the key property of exhibiting a purely Tellegen response, without gyroelectric, gyromagnetic, and other bianisotropic effects.
To avoid limitations imposed by specific nanofabrication technologies, we consider idealized nanocones, as depicted in the inset of Fig. 1(a), for both material systems in Figs. 4(a) and (d). We also assume that the nanocones remain in an ideal single-domain state in the absence of an external magnetic field. Using full-wave simulations, we determine the collective polarizabilities $\alpha_{\rm ge}$, $\alpha_{\rm gm}$, and $\alpha_{\rm te}$ for the Tellegen membrane (Supplementary Section 3). As shown in Fig. 4(b), the membrane exhibits only the Tellegen response. Indeed, due to the symmetry of the unit cell, the gyroelectric and gyromagnetic effects are nearly completely suppressed [8]. The Tellegen polarizability $\alpha_{\rm te}$ in the membrane is approximately twice larger than in the single metasurface (see Fig. 1(d)) since the unit cell includes two cones instead of one. In terms of observable effects, this membrane demonstrates the so-called isotropic Kerr effect [10], implying that, regardless of the illumination direction, the Kerr magneto-optical parameters exhibit the same sign, i.e., $\theta^{\rm f}=\theta^{\rm b}$ and $\epsilon^{\rm f}=\epsilon^{\rm b}$ (here, the Kerr rotation direction is defined with respect to the illumination direction). These parameters are plotted in Fig. 4(c). This property stands in sharp contrast to conventional gyroelectric or gyromagnetic metasurfaces, where the Kerr parameters have opposite signs for opposite illumination directions. Moreover, the Tellegen membrane exhibits no Faraday effect. For comparison, in Supplementary Section 8, we consider the same membrane geometry but where the magnetization in nanocones on both sides is the same. In that case, the symmetry results in a zero Tellegen effect.
To evaluate the macroscopic effective-medium parameters of the purely Tellegen colloid shown in Fig. 4(d), we apply the Maxwell-Garnett mixing rule [15, Sec. 7.2.2]. Here, we assume the host medium to be water, however, other dielectric media would yield comparable results, differing mainly in the effective permittivity of the colloid. The volumetric concentration of the meta-atoms is set to $N=3.47\times 10^{20}\leavevmode\nobreak\ {\rm m}^{-3}$, corresponding to an experimentally feasible volume fraction of the cones $N_{V}=0.1$. Figure 4(e) presents the effective bulk calculated Tellegen parameter, $\chi_{\rm eff}$, exhibiting a Lorentzian resonance spectral shape. At resonance near 850 nm, $\chi_{\rm eff}$ reaches approximately $10^{-3}$, which is at least 100 times larger than values observed in natural materials such as chromium oxide [47]. When using the experimentally calculated Tellegen polarizability from Fig. 3(i) in our theoretical calculations, we obtain a slightly smaller value of the Tellegen bulk parameter, $\chi_{\rm eff}=8\times 10^{-4}$, at the resonance. Owing to its isotropic response, the Tellegen colloid exhibits identical magneto-optical behavior regardless of the illumination direction. Figure 4(f) plots these “isotropic” Kerr parameters for the colloid, showing qualitative similarity to those of the membrane in Fig. 4(c), as both systems exhibit a purely Tellegen response.
In conclusion, we have demonstrated the first experimental realization of the optical Tellegen effect in artificial metamaterial system, with the strength of around 100 times greater than that of natural materials. The Tellegen metasurface was produced using conventional materials and a well-established, highly-scalable nanofabrication, ensuring practicality for diverse applications. Since the presented optical metasurface also exhibits gyroelectric and gyromagnetic effects, we introduced an approach for independently extracting each of the three nonreciprocal effects. This technique is expected to be instrumental in the magneto-optical characterization of other nonreciprocal optical metasurfaces, particularly where symmetry permits multiple nonreciprocal effects. Experimental measurements, including those of magneto-optical Kerr parameters and nonreciprocal polarizabilities, demonstrate strong agreement with full-wave simulations, confirming the accuracy and reliability of the proposed method.
The inherent scalability of the chosen bottom-up nanofabrication approach facilitates the straightforward extension of the proposed nonreciprocal optical Tellegen metasurface to other metamaterial systems, including bulk Tellegen metamaterials and colloids. By tailoring the geometry and composition of meta-atoms, the resonant wavelength can be precisely tuned across the optical spectrum. Further, opening-up the ability to modulate the Tellegen metamaterial in space and/or time has profound implications for generating and analyzing effective axion fields— which are inherently nonuniform [54, 55, 56]— and other related phenomena [2, 3, 4, 57, 58].
Materials and methods
Simulation
The full-wave simulation characterization of the metasurfaces was performed using COMSOL Multiphysics software with a frequency-domain solver. Periodic boundary conditions were applied to simulate an infinite domain around the structure, with the polarization of the incident light defined at the incident port. The complex reflection spectra were then calculated based on the scattering parameters. We used the material dispersion properties of the alumina, aluminum, cobalt, and amorphous silicon from measurements reported in [59], [60], [61, 62], and [63], respectively. To account for the inevitable oxidation of the bare aluminum layer in the first fabricated metasurface sample (which has a target alumina thickness of $h_{1}=0$ nm), we incorporated an alumina oxidation layer with a thickness of 9 nm in our full-wave simulations used to generate the results in Fig. 3(d).
Sample fabrication
To achieve large-scale fabrication of optical Tellegen metasurfaces, we employed hole-mask colloidal lithography [64]. Microscope slide glass substrates (VWR International) with dimensions of 24 mm × 24 mm × 0.4 mm were used. The substrates were cleaned through an ultrasonication process in acetone. Subsequently, a 5 nm titanium (Ti) adhesion layer and a 100 nm aluminum (Al) layer were deposited using electron-beam physical vapor deposition (EB-PVD). Atomic layer deposition (ALD) was utilized to deposit aluminum oxide (Al2O3) films with thicknesses of 60 nm and 120 nm for the second and third metasurfaces, respectively. A polymethyl methacrylate (PMMA) layer (950PMMA, 4% solution in anisole, MW = 950,000, MicroChem) with a thickness of 250 nm was spin-coated on all metasurfaces and soft-baked at 130°C for 3 minutes on a hot plate. Following the baking process, the PMMA layer was subjected to oxygen plasma treatment for 15 seconds to enhance its hydrophilicity. The PMMA-coated substrate was then functionalized with polydiallyldimethylammonium chloride (PDDA, 0.2% solution in deionized (DI) water), followed by rinsing with DI water and drying with a nitrogen (N2) gun. Negatively charged polystyrene beads (100 nm, sulfate latex, 0.2% suspension in DI water, Invitrogen) were deposited onto the substrate via pipetting and left in contact for 120 seconds. After careful rinsing with DI water, the samples were dried using an N2 gun at normal flow incidence to prevent bead aggregation. A 10 nm chromium (Cr) layer was subsequently deposited, followed by tape-stripping of the polystyrene spheres (SWT-10 tape, Nitto Scandinavia AB). Reactive ion etching (RIE) with oxygen plasma was then performed to etch through the entire PMMA layer, forming an evaporation mask. The silicon (Si) and cobalt (Co) layers were deposited using EB-PVD. During material deposition, a progressive reduction in the nanohole diameter occurred as a result of the accumulation of material along the hole edges, leading to a decrease in the diameter of the nanodiscs and the formation of conical meta-atoms. The final step involved the removal of the PMMA layer via acetone lift-off.
Magneto-optical characterization
The magneto-optical Kerr effect (MOKE) of the samples was characterized using an NKT SuperK EXW-12 supercontinuum laser connected to an acousto-optical tunable filter (AOTF), providing a wavelength range of 500–1400 nm. The AOTF employed in our setup has two outputs: VIS/NIR (500–900 nm) and NIR2 (800–1400 nm). A focusing doublet, Glan-Thompson polarizer, and silver D-shaped mirror were used to direct the light into the magnet poles and focus it onto the samples, which were positioned between the poles of a 3470 GMW electromagnet (Fig. 3(c)). The measurements were performed at normal incidence. The reflected beam from the samples passed sequentially through a photoelastic modulator I/FS50 Hinds Instruments (PEM), a half-wave plate Thorlabs FR600HM (HWP), and a beam splitter Thorlabs WPA10 (BS), before being collected by a balanced photodetector Thorlabs PDB250A2 (BPD). In the MOKE setup, the combination of the photoelastic modulator with two lock-in amplifiers operating at 50 kHz (Signaloc 2100, Hinds Instruments) and 100 kHz (SR830 DSP, Stanford Research Systems), along with a balance detector, enabled the simultaneous measurement of Kerr rotation ($\theta$) and Kerr ellipticity ($\varepsilon$).
Data Availability
The authors declare that the data supporting the findings of this study are available within the paper, in the Supplementary file, and are available from the corresponding authors upon request.
Acknowledgment
S.S.J. and V.A. acknowledge the Academy of Finland (Project No. 356797). V.A. acknowledges the PREIN Flagship (Project No. 320167) and the Finnish Foundation for Technology Promotion. S. F. acknowledges a MURI project from the U. S. Army Research Office (Grant No. W911NF-24-1-0224). I.F., R.C., and A.D. acknowledge the Swedish Research Council for Sustainable Development (Formas) (Project No. 2021-01390), Thuréus Forskarhem och Naturminne Foundation, and the Swedish Research Council (VR) (Project No. 2024-05025). N.K. acknowledges financial support from the Magnus Ehrnrooth Foundation. We acknowledge the Micronova Nanofabrication Centre, supported by Aalto University, for providing resources for this work. We also thank Julius Hohlfeld and Junyoung Hyun for their assistance with the measurement setup.
Author contributions
V.A. conceived the idea and determined the optimal topology of the meta-atom. S.S.J. and V.A. conducted the theoretical calculations. S.S.J. and I.F. performed the full-wave simulations of empirical samples. I.F. and S.S.J. were responsible for the fabrication. S.S.J. and R.C. conducted the experimental measurements and N.K. provided assistance and support during the measurement process. S.S.J., V.A, and I.F. characterized the experimental and simulated results. S.v.D., S.F., A.D., and V.A. supervised the work. S.S.J. drafted the manuscript, with all authors contributing to its review and revision. All authors contributed to the discussions of the results.
Competing interests
The authors declare no competing interests.